# A Copula Based Supervised Filter for Feature Selection in Machine Learning Driven Diabetes Risk Prediction

### v4

@@ Abstract
Effective feature selection is critical for building robust and interpretable predictive models, particularly in medical applications where identifying risk factors in the most extreme patient strata is essential. Traditional methods often focus on average associations, potentially overlooking predictors whose importance is concentrated in the tails of the data distribution. In this study, we introduce a novel, computationally efficient supervised filter that leverages the Gumbel copula’s upper-tail dependence coefficient ($\lambda_{U}$) to rank features by their tendency to be simultaneously extreme with the positive class. We evaluated this method against four standard baselines (Mutual Information, mRMR, ReliefF, and L1/Elastic-Net) across four classifiers on two diabetes datasets: a large-scale public health survey (CDC, N=253,680N{=}253{,}680) and a classic clinical benchmark (PIMA, N=768N{=}768). Our analysis included comprehensive statistical tests, permutation importance, and robustness checks. On the CDC dataset, our method was the fastest selector and reduced the feature space by $\approx$52%. While this resulted in a minimal but statistically significant performance trade-off compared to using all 21 features, our filter significantly outperformed standard filters (Mutual Information, mRMR) and was statistically indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our method’s ranking produced the numerically highest ROC-AUC, despite paired DeLong tests showing no statistically significant differences versus strong baselines. PIMA thus serves as a ranking-only sanity check that our upper-tail criterion behaves sensibly in a low-dimensional clinical setting. Across both datasets, the Gumbel-$\lambda_{U}$ selector consistently identified clinically coherent and impactful predictors. We conclude that feature selection via upper-tail dependence is a powerful, efficient, and interpretable tool for developing risk models in public health and clinical medicine.
@@ Introduction
We introduce and validate a supervised filter for feature selection based on the upper-tail dependence coefficient $\lambda_{U}$ of the Gumbel copula [7]. Each predictor is ranked by its $\lambda_{U}(X,Y)$ with the diabetes label, prioritizing features whose highest values co-occur with positive outcomes. To our knowledge, this is the first study to operationalize a copula tail-dependence coefficient as a direct, standalone criterion for supervised feature selection in clinical risk prediction.
We validate on two public datasets. On the large CDC survey, our method demonstrates a powerful trade-off between parsimony and performance. By reducing the feature set by 52% (from 21 to 10), our Gumbel-selected model achieves a competitive ROC-AUC of 0.823 and is the fastest selector. Most importantly, it proves its efficacy as a top-tier filter by statistically outperforming standard filter methods like Mutual Information and mRMR, while performing on par with the strong ReliefF baseline. To demonstrate external validity beyond a large, survey-style cohort (CDC), we also evaluate on the PIMA Indians Diabetes dataset, a small clinical benchmark with eight predictors. Because PIMA contains all and only eight covariates, this experiment isolates feature ranking rather than dimensionality reduction: all methods use the same variables. The Gumbel ranking combined with a Random Forest attains the highest ROC-AUC of 0.867, with paired DeLong tests showing no significant differences across selectors (DeLong $p>0.05$). As such, PIMA functions as a ranking-only sanity check that our upper-tail criterion yields clinically coherent orderings without harming discrimination. Overall, our results show that focusing on the extremes can create models that are simple, easy to understand, and effective for clinical and public health use.
@@
Importantly, tail-dependence coefficients, which directly measure the probability of joint extremes, have not yet been used as the main supervised ranking criterion in feature selection. As seen from the literature, a Gaussian copula cannot capture tail dependence and have called for methods that use tail-dependent families. Our work addresses this need. By using the Gumbel copula’s upper-tail coefficient $\lambda_{U}$ [11, 7], we shift the selection focus from overall association to joint extremal behavior. This approach provides a simple, interpretable, and computationally efficient supervised filter that (i) identifies clinically meaningful predictors linked to high-risk cases, (ii) matches top performance on both large survey and clinical datasets, and (iii) remains fast compared to standard baselines like MI, mRMR, and ReliefF, making it practical for real-world use.
@@
Scope. Tail-dependence coefficients are defined for any bivariate distribution and, via Sklar’s theorem, for any copula family.
@@
We use a property-first approach and require a dependence measure with nonzero $\lambda_{U}$. The Gumbel copula meets this requirement, since $\lambda_{U}>0$ and $\lambda_{L}=0$. In contrast, the Clayton copula only has lower-tail dependence ($\lambda_{U}=0$), and the Gaussian and Frank copulas are tail-independent ($\lambda_{U}=\lambda_{L}=0$). Even if a goodness-of-fit test favored those other families, they would not address our question about upper-tail synchrony. For this reason, we use the Gumbel mapping as a monotone transform of $\tau$ to rank features by upper-tail dependence. We are not claiming a full generative copula fit.
@@
We standardize inputs only where the downstream method requires distance or coefficient comparability:(i) LR is trained on standardized predictors (scaler fit on train only and applied to val/test) to stabilize the optimizer and penalty. (ii) ReliefF relies on Euclidean nearest neighbors, so we z-score each feature on the training split only before its hit/miss computations and apply the learned mean/SD to val/test. By contrast, (iii) tree ensembles (RF/GB/XGB) are scale-invariant and use raw inputs; and (iv) the MI and mRMR filters are used without scaling: MI here is univariate $I(X_{j};Y)$, so linear rescaling of $X_{j}$ does not change the information content, and the mRMR redundancy term uses Pearson correlation, which is itself scale-invariant. All scalers are train-only to avoid leakage.
@@
For each feature $X_{j}$, we compute the Gumbel upper–tail dependence score $\lambda_{U}(X_{j},Y)$ from rank-based pseudo-observations and rank features by $\lambda_{U}$. We say a feature shows a positive upper-tail association when the 95% bootstrap confidence interval for $\lambda_{U}$ lies above 0. We keep the top-$k$ features and train models on that set. To judge usefulness, we compare these models with baselines (MI, mRMR, ReliefF, and L1/Elastic-Net) on the same test split, using ROC-AUC as the primary metric (with PR–AUC, accuracy, recall, and F1 as secondary metrics). For formal paired comparisons we use DeLong’s test for ROC-AUC and McNemar’s test for error profiles.
@@
We calculate Kendall’s $\tau$ for each feature against the label. Then, we map $\tau$ to the Gumbel copula parameter with the formula $\hat{\theta}=1/(1-\tau)$ whenever $\tau>0$. After that, we convert $\hat{\theta}$ into the upper-tail dependence coefficient $\lambda_{U}=2-2^{1/\hat{\theta}}$. We rank the features based on $\lambda_{U}$. To measure uncertainty in these scores, we bootstrap the training data 1,000 times with replacement. Each time, we recompute $\lambda_{U}$ and report the bootstrap mean along with the 95% percentile interval. .
@@ Results
Table 3: Predictive performance on the CDC dataset test set for each feature selection method and classifier. The highest ROC-AUC within each feature set is bolded.
Set Model Accuracy Precision Recall F1-Score ROC-AUC
All GB 0.804 0.854 0.804 0.822 0.827
LR 0.791 0.852 0.791 0.813 0.820
RF 0.779 0.847 0.779 0.804 0.801
XGB 0.809 0.853 0.809 0.826 0.827
Gumbel GB 0.813 0.850 0.813 0.827 0.823
LR 0.797 0.849 0.797 0.817 0.816
RF 0.732 0.826 0.732 0.766 0.741
XGB 0.810 0.851 0.810 0.826 0.822
MI GB 0.821 0.848 0.821 0.832 0.821
LR 0.793 0.849 0.793 0.814 0.815
RF 0.733 0.833 0.733 0.768 0.755
XGB 0.814 0.849 0.814 0.828 0.820
mRMR GB 0.781 0.843 0.781 0.804 0.794
LR 0.763 0.843 0.763 0.791 0.789
RF 0.745 0.839 0.745 0.777 0.763
XGB 0.781 0.842 0.781 0.804 0.792
ReliefF GB 0.805 0.851 0.805 0.822 0.822
LR 0.791 0.850 0.791 0.813 0.815
RF 0.748 0.835 0.748 0.779 0.766
XGB 0.814 0.849 0.814 0.828 0.821
L1EN GB 0.801 0.855 0.801 0.820 0.826
LR 0.804 0.850 0.804 0.822 0.819
RF 0.733 0.833 0.733 0.768 0.753
XGB 0.809 0.852 0.809 0.825 0.825
@@
Using all 21 features gives the best overall performance, which is common in datasets with relatively few features. However, our main focus is to assess how well the selection methods work. The results in Table 3 show a clear ranking. Our Gumbel-$\lambda_{U}$ method and the L1/Elastic-Net (L1EN) baseline consistently perform at the top, with strong and nearly identical results across all classifiers. For example, with the Gradient Boosting (GB) model, Gumbel-$\lambda_{U}$ achieved a ROC-AUC of 0.823, which is very close to L1EN’s 0.826. This shows that our copula-based filter method is highly competitive with a leading embedded selector. When using Gumbel-$\lambda_{U}$ selected features, the Gradient Boosting (GB) model achieved the highest ROC-AUC of 0.823, establishing it as the best-performing model for our proposed feature set.
@@
The Gumbel-$\lambda_{U}$ feature set shows a clear advantage over several other filter methods. Its ROC-AUC of 0.823 is numerically higher than the scores for Mutual Information (MI) at 0.821 and mRMR at 0.794. As the formal statistical tests in the next section will confirm, our method is not only competitive but statistically superior to these common filters, indicating that its focus on upper-tail dependence captures a more relevant predictive signal.
@@
To provide a visual comparison of these results, Figure 1 displays the superimposed ROC curves for the top-performing Gradient Boosting model across each feature set. The plot clearly shows the curves for our Gumbel-$\lambda_{U}$ method, MI, L1EN, ReliefF and the All Features baseline clustered tightly together, visually confirming their nearly identical and superior performance. In contrast, the curve for mRMR falls visibly below this top group, consistent with their lower AUC score reported in Table 3. This visualization reinforces the conclusion that our proposed method effectively identifies a feature set that is highly competitive with the best-performing baselines.
@@
Table 8: Predictive performance on the PIMA dataset test set for each feature ranking and classifier.
Set Model Accuracy Precision Recall F1-Score ROC-AUC
All GB 0.773 0.806 0.773 0.778 0.860
LR 0.773 0.768 0.773 0.766 0.862
RF 0.766 0.792 0.766 0.771 0.861
XGB 0.779 0.786 0.779 0.782 0.840
Gumbel GB 0.682 0.790 0.682 0.685 0.858
LR 0.773 0.768 0.773 0.766 0.862
RF 0.753 0.801 0.753 0.759 0.867
XGB 0.760 0.804 0.760 0.765 0.824
MI GB 0.773 0.806 0.773 0.778 0.859
LR 0.773 0.768 0.773 0.766 0.862
RF 0.812 0.817 0.812 0.813 0.860
XGB 0.747 0.776 0.747 0.752 0.842
mRMR GB 0.760 0.799 0.760 0.765 0.863
LR 0.773 0.768 0.773 0.766 0.862
RF 0.760 0.799 0.760 0.765 0.863
XGB 0.753 0.784 0.753 0.759 0.839
ReliefF GB 0.779 0.815 0.779 0.784 0.866
LR 0.773 0.768 0.773 0.766 0.862
RF 0.812 0.813 0.812 0.812 0.860
XGB 0.799 0.804 0.799 0.801 0.835
L1EN GB 0.760 0.817 0.760 0.765 0.864
LR 0.773 0.768 0.773 0.766 0.862
RF 0.792 0.809 0.792 0.796 0.861
XGB 0.753 0.775 0.753 0.758 0.833
On this dataset, our Gumbel-$\lambda_{U}$ ranking yielded the numerically highest ROC-AUC. The top value of 0.867 was obtained by pairing the Gumbel-ranked features with a Random Forest (RF) classifier. This edges strong alternatives (e.g., ReliefF: 0.866 with GB; L1EN: 0.864 with GB). In a low-dimensional clinical setting where feature removal is not the primary goal, this indicates that a tail-aware ranking is competitive while remaining interpretable, rather than conclusively superior.
@@
On CDC (21 features), upper-tail filtering reduces the feature set by $\sim$52%. This introduces a minor but statistically significant performance trade-off against using the full feature set, but our method’s key advantage is highlighted by its statistically superior performance compared to other filter baselines, demonstrating its efficiency and ranking power. On PIMA (8 features), where selection is not applicable, our criterion yields clinically coherent rankings and maintains discrimination (no significant AUC differences), i.e., it does no harm in low-dimensional clinical settings. Together, these results support the method’s robustness, interpretability, and practicality across dataset types.
@@
In this study, we developed a new supervised filter method for feature selection using the Gumbel copula’s upper-tail dependence coefficient, $\lambda_{U}$. By focusing on how extreme values in predictors and the target label occur together, our method provides a fresh way to find features that matter most in high-risk situations. Our thorough evaluation, shaped by early reviewer feedback, shows that this approach is statistically solid, effective, and efficient in real-world use.
We tested our method on two different diabetes datasets and found strong support for its usefulness. On the large CDC public health dataset, our Gumbel-$\lambda_{U}$ selector was competitive and was the fastest method, running much quicker than other approaches. It also reduced the number of features by about $\approx$ 52% (from 21 to 10), and on this parsimonious set, it proved its value as a top-tier filter by delivering performance that was statistically superior to standard filters like Mutual Information and mRMR and indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our ranking achieved the numerically highest ROC-AUC while remaining statistically indistinguishable from strong baselines, functioning as a ranking-only clinical sanity check. Although it was not the fastest on this smaller dataset, its speed was still close to other standard filters, showing that its strong performance did not come at a high computational cost. In both cases, our method consistently found clinically important predictors, showing it can identify meaningful and understandable signals.
@@
The idea of finding features that are extreme alongside high-risk outcomes can be used in many areas of biomedical research, especially where there is a lot of data. For example, in genomics and transcriptomics, this approach can help identify genes that are highly active in aggressive cancers. In proteomics and metabolomics, it can find molecules that are most abundant in severe disease or in response to treatment. This method can also improve clinical trial analysis by spotting unusual lab values that predict which patients might benefit most or be at higher risk, making it easier to focus on the right groups. In neuroimaging, it provides a new way to look at fMRI or PET scans by highlighting brain regions with unusually high activity linked to certain neurological disorders, adding to what we learn from average-based analyses. All of these uses aim to build better personalized risk models, helping to identify patients who are most at risk so they can get early and targeted care.
@@
In summary, using copula-based upper-tail dependence offers a strong, efficient, and understandable way to select features. By finding predictors that appear with outcomes in the highest-risk groups and performing as well as or better than strong baselines on both large and small clinical datasets, this method is a useful tool for machine learning, public health, and clinical medicine. The suggested extensions above show clear ways to strengthen its theory, handle real-world data challenges, and expand its impact in biomedical fields.
@@
The CDC Diabetes Health Indicators dataset is available from the UCI Machine Learning Repository [2]. The PIMA Indians Diabetes dataset originates from NIDDK and Smith et al.[41]; a curated copy used for this work was accessed via Kaggle.
@@
All code to reproduce the analyses is available at github.com/agnivibes/copula-feature-selection.

### v5

@@ Abstract
Effective feature selection is critical for building robust and interpretable predictive models, particularly in medical applications where identifying risk factors in the most extreme patient strata is essential. Traditional methods often focus on average associations, potentially overlooking predictors whose importance is concentrated in the tails of the data distribution. In this study, we introduce a novel, computationally efficient supervised filter that leverages a Gumbel copula implied upper-tail concordance score ($\lambda_{U}$, a monotone transformation of Kendall’s $\tau$) to rank features by their tendency to be simultaneously extreme with the positive class. We evaluated this method against four standard baselines (Mutual Information, mRMR, ReliefF, and L1/Elastic-Net) across four classifiers on two diabetes datasets: a large-scale public health survey (CDC, N=253,680N{=}253{,}680) and a classic clinical benchmark (PIMA, N=768N{=}768). Our analysis included comprehensive statistical tests, permutation importance, and robustness checks. On the CDC dataset, our method was the fastest selector and reduced the feature space by $\approx$52%. While this resulted in a minimal but statistically significant performance trade-off compared to using all 21 features, our filter significantly outperformed standard filters (Mutual Information, mRMR) and was statistically indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our method’s ranking produced the numerically highest ROC-AUC, despite paired DeLong tests showing no statistically significant differences versus strong baselines. PIMA thus serves as a ranking-only sanity check that our upper-tail criterion behaves sensibly in a low-dimensional clinical setting. Across both datasets, the Gumbel-$\lambda_{U}$ selector consistently identified clinically coherent and impactful predictors. We conclude that feature selection via upper-tail dependence is an efficient and interpretable screening approach that can complement standard feature-selection baselines in public health and clinical risk prediction.
@@ Introduction
We introduce and validate a supervised filter for feature selection based on the Gumbel copula’s [7] upper-tail dependence coefficient $\lambda_{U}$, used here as a tail-sensitive concordance score derived from rank-based dependence. Each predictor is ranked by its $\lambda_{U}(X,Y)$ with the diabetes label, prioritizing features whose highest values co-occur with positive outcomes. To our knowledge, this is the first study to operationalize a copula tail-dependence coefficient as a direct, standalone criterion for supervised feature selection in clinical risk prediction.
We validate on two public datasets. On the large CDC survey, our method demonstrates a favorable trade-off between parsimony and performance. By reducing the feature set by 52% (from 21 to 10), our Gumbel-selected model achieves a competitive ROC-AUC of 0.823 and is the fastest selector. Most importantly, it demonstrates its utility as a competitive filter by achieving statistically higher AUC scores than standard filter methods like Mutual Information and mRMR, while performing on par with the strong ReliefF baseline. To demonstrate external validity beyond a large, survey-style cohort (CDC), we also evaluate on the PIMA Indians Diabetes dataset, a small clinical benchmark with eight predictors. Because PIMA contains all and only eight covariates, this experiment isolates feature ranking rather than dimensionality reduction: all methods use the same variables. The Gumbel ranking combined with a Random Forest attains the highest ROC-AUC of 0.867, with paired DeLong tests showing no significant differences across selectors (DeLong $p>0.05$). As such, PIMA functions as a ranking-only sanity check that our upper-tail criterion yields clinically coherent orderings without harming discrimination. Overall, our results show that focusing on the extremes can create models that are simple, easy to understand, and effective for clinical and public health use.
@@
Importantly, tail-dependence coefficients, which directly measure the probability of joint extremes, have not yet been used as the main supervised ranking criterion in feature selection. As seen from the literature, a Gaussian copula cannot capture tail dependence and have called for methods that use tail-dependent families. Our work addresses this need. By using the Gumbel copula’s upper-tail coefficient $\lambda_{U}$ [11, 7], we shift the selection focus from overall association to joint extremal behavior. This approach provides a simple, interpretable, and computationally efficient supervised filter that (i) identifies clinically meaningful predictors linked to high-risk cases, (ii) achieves performance comparable to strong baselines on both the large survey and clinical benchmark datasets, and (iii) remains fast compared to standard baselines like MI, mRMR, and ReliefF, making it practical for real-world use.
@@
Scope. Tail-dependence concepts are most naturally defined for continuous margins; in this paper, we use the Gumbel mapping as a rank-based, tail-sensitive concordance score (a monotone transformation of Kendall’s $\tau$) for feature ranking, rather than as a generative copula tail parameter. Throughout the paper, references to $\lambda_{U}$ (upper-tail dependence) are to this Gumbel-implied, rank-based concordance score derived from Kendall’s $\tau$, used for ranking rather than for generative copula modeling.
@@
We use a property-first approach and require a dependence measure with nonzero $\lambda_{U}$. The Gumbel copula meets this requirement, since $\lambda_{U}>0$ and $\lambda_{L}=0$. In contrast, the Clayton copula only has lower-tail dependence ($\lambda_{U}=0$), and the Gaussian and Frank copulas are tail-independent ($\lambda_{U}=\lambda_{L}=0$). Even if a goodness-of-fit test favored those other families, they would not address our question about upper-tail synchrony. For this reason, we use the Gumbel mapping as a monotone transform of $\tau$ to rank features by upper-tail dependence. We are not claiming a full generative copula fit. While the classical upper-tail dependence coefficient is defined asymptotically for continuous margins, here it is used as a monotone transformation of Kendall’s $\tau$ to construct a tail-sensitive concordance score for ranking; we do not claim estimation of asymptotic tail probabilities or a fully specified generative copula model.
@@
We standardize inputs only where the downstream method requires distance or coefficient comparability:(i) LR is trained on standardized predictors (scaler fit on train only and applied to val/test) to stabilize the optimizer and penalty. (ii) ReliefF relies on Euclidean nearest neighbors, so we z-score each feature on the training split only before its hit/miss computations and apply the learned mean/SD to val/test. (iii) L1EN also operates on features that have been scaled to have a zero mean and unit variance. By contrast, (iv) tree ensembles (RF/GB/XGB) are scale-invariant and use raw inputs; and (v) the MI and mRMR filters are used without scaling: MI here is univariate $I(X_{j};Y)$, so linear rescaling of $X_{j}$ does not change the information content, and the mRMR redundancy term uses Pearson correlation, which is itself scale-invariant. All scalers are train-only to avoid leakage.
@@
For each feature $X_{j}$, we compute the Gumbel upper–tail dependence score $\lambda_{U}(X_{j},Y)$ from rank-based pseudo-observations and rank features by $\lambda_{U}$. We say a feature shows a positive upper-tail association when the 95% bootstrap percentile confidence interval for $\lambda_{U}(X_{j},Y)$, computed from $B=1000$ bootstrap resamples of the training split (with replacement), lies strictly above 0. We keep the top-$k$ features and train models on that set. To judge usefulness, we compare these models with baselines (MI, mRMR, ReliefF, and L1/Elastic-Net) on the same test split, using ROC-AUC as the primary metric (with PR–AUC, accuracy, recall, and F1 as secondary metrics). For ROC-AUC, we also report 95% bootstrap percentile confidence intervals on the test set using $B=1000$ resamples of test indices. For formal paired comparisons we use DeLong’s test for ROC-AUC and McNemar’s test for error profiles.
@@
We calculate Kendall’s $\tau$ for each feature against the label. Then, we map $\tau$ to the Gumbel copula parameter with the formula $\hat{\theta}=1/(1-\tau)$ whenever $\tau>0$. After that, we convert $\hat{\theta}$ into the upper-tail dependence coefficient $\lambda_{U}=2-2^{1/\hat{\theta}}$. We rank the features based on $\lambda_{U}$. To quantify uncertainty in the tail-dependence scores, we bootstrap the training split $B=1000$ times with replacement, recompute $\lambda_{U}$ for each resample, and report the bootstrap mean and the 2.5th–97.5th percentile interval.
@@ Results
Table 3: Predictive performance on the CDC dataset test set for each feature selection method and classifier. The highest ROC-AUC within each feature set is bolded. ROC-AUC is reported with 95% bootstrap confidence intervals.
Set Model Accuracy Precision Recall F1-Score ROC-AUC (95% CI)
All GB 0.804 0.854 0.804 0.822 0.827 [0.823, 0.832]
LR 0.791 0.852 0.791 0.813 0.820 [0.815, 0.825]
RF 0.779 0.847 0.779 0.804 0.801 [0.796, 0.806]
XGB 0.809 0.853 0.809 0.826 0.827 [0.823, 0.832]
Gumbel GB 0.813 0.850 0.813 0.827 0.823 [0.819, 0.828]
LR 0.797 0.849 0.797 0.817 0.816 [0.812, 0.821]
RF 0.732 0.826 0.732 0.766 0.741 [0.735, 0.746]
XGB 0.810 0.851 0.810 0.826 0.822 [0.817, 0.827]
MI GB 0.821 0.848 0.821 0.832 0.821 [0.817, 0.826]
LR 0.793 0.849 0.793 0.814 0.815 [0.810, 0.820]
RF 0.733 0.833 0.733 0.768 0.755 [0.748, 0.760]
XGB 0.814 0.849 0.814 0.828 0.820 [0.815, 0.825]
mRMR GB 0.781 0.843 0.781 0.804 0.794 [0.788, 0.799]
LR 0.763 0.843 0.763 0.791 0.789 [0.784, 0.795]
RF 0.745 0.839 0.745 0.777 0.763 [0.757, 0.769]
XGB 0.781 0.842 0.781 0.804 0.792 [0.787, 0.797]
ReliefF GB 0.805 0.851 0.805 0.822 0.822 [0.818, 0.827]
LR 0.791 0.850 0.791 0.813 0.815 [0.810, 0.820]
RF 0.748 0.835 0.748 0.779 0.766 [0.760, 0.771]
XGB 0.814 0.849 0.814 0.828 0.821 [0.816, 0.826]
L1EN GB 0.801 0.855 0.801 0.820 0.826 [0.822, 0.831]
LR 0.804 0.850 0.804 0.822 0.819 [0.815, 0.824]
RF 0.733 0.833 0.733 0.768 0.753 [0.747, 0.759]
XGB 0.809 0.852 0.809 0.825 0.825 [0.820, 0.830]
We report ROC-AUC with 95% bootstrap confidence intervals ($B=1000$ resamples) to quantify uncertainty under resampling and to complement paired significance tests.
@@
Using all 21 features gives the best overall performance, which is common in datasets with a moderate number of predictors. Moreover, the All, Gumbel-$\lambda_{U}$, and L1EN configurations yield closely matched ROC-AUC values, and their 95% bootstrap confidence intervals frequently overlap, indicating that the observed differences are small relative to resampling variability. Our main focus, however, is to assess how well the selection methods perform. The results in Table 3 show a clear ranking: our Gumbel-$\lambda_{U}$ method and the L1/Elastic-Net (L1EN) baseline consistently perform among the top methods across classifiers, with strong and nearly identical results. For example, with Gradient Boosting (GB), Gumbel-$\lambda_{U}$ attains a ROC-AUC of 0.823, which is close to L1EN’s 0.826. Within the Gumbel-$\lambda_{U}$ feature set, GB achieves the highest ROC-AUC (0.823), making it the best-performing classifier for our proposed selected features on the CDC dataset.
@@
The Gumbel-$\lambda_{U}$ feature set shows a clear advantage over several other filter methods. Its ROC-AUC of 0.823 is numerically higher than the scores for Mutual Information (MI) at 0.821 and mRMR at 0.794. As the formal statistical tests in the next section will confirm, our method is not only competitive but statistically higher ROC-AUC than MI and mRMR on this split, suggesting that the upper-tail criterion can be a useful screening signal in this setting.
@@
To provide a visual comparison of these results, Figure 1 displays the superimposed ROC curves for the top-performing Gradient Boosting model across each feature set. The plot clearly shows the curves for our Gumbel-$\lambda_{U}$ method, MI, L1EN, ReliefF and the All Features baseline clustered tightly together, visually confirming their very similar performance.. In contrast, the curve for mRMR falls visibly below this top group, consistent with their lower AUC score reported in Table 3. This visualization reinforces the conclusion that our proposed method effectively identifies a feature set that is highly competitive with the best-performing baselines.
@@
Comment on CDC feature type and model behavior.
The CDC dataset contains a large number of binary/discrete predictors. In such settings, a linear model can perform competitively because much of the apparent nonlinearity is effectively represented through indicator coding, which helps explain why logistic regression can match or exceed tree-based models in our benchmarks. As a result, while the CDC experiment provides strong evidence that the proposed upper-tail screening is effective and efficient, further evaluation on datasets with richer continuous predictors and stronger interaction-driven nonlinearities is a natural direction for future work.
@@
Table 8: Predictive performance on the PIMA dataset test set for each feature ranking and classifier. We report ROC-AUC with 95% bootstrap confidence intervals. The highest ROC-AUC within each feature set is bolded.
Set Model Accuracy Precision Recall F1 ROC-AUC (95% CI)
All GB 0.773 0.806 0.773 0.778 0.860 [0.791, 0.916]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.766 0.792 0.766 0.771 0.861 [0.798, 0.918]
XGB 0.779 0.786 0.779 0.782 0.840 [0.767, 0.899]
Gumbel GB 0.682 0.790 0.682 0.685 0.858 [0.789, 0.914]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.753 0.801 0.753 0.759 0.867 [0.804, 0.920]
XGB 0.760 0.804 0.760 0.765 0.824 [0.748, 0.885]
MI GB 0.773 0.806 0.773 0.778 0.859 [0.789, 0.914]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.812 0.817 0.812 0.813 0.859 [0.794, 0.912]
XGB 0.747 0.776 0.747 0.752 0.842 [0.769, 0.902]
mRMR GB 0.760 0.799 0.760 0.765 0.863 [0.795, 0.919]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.760 0.799 0.760 0.765 0.863 [0.796, 0.917]
XGB 0.753 0.784 0.753 0.759 0.839 [0.768, 0.898]
ReliefF GB 0.779 0.815 0.779 0.784 0.866 [0.799, 0.920]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.812 0.813 0.812 0.812 0.860 [0.795, 0.915]
XGB 0.799 0.804 0.799 0.801 0.835 [0.765, 0.896]
L1EN GB 0.760 0.817 0.760 0.765 0.864 [0.799, 0.918]
LR 0.773 0.768 0.773 0.766 0.862 [0.796, 0.913]
RF 0.792 0.809 0.792 0.796 0.861 [0.800, 0.913]
XGB 0.753 0.775 0.753 0.758 0.833 [0.763, 0.894]
We report ROC-AUC with 95% bootstrap confidence intervals ($B=1000$ resamples) to quantify uncertainty under resampling and to complement paired significance tests.
On PIMA (8 predictors), dimensionality reduction is not the goal, so this experiment primarily serves as a ranking sanity check in a low-dimensional clinical setting. In Table 8, the best-performing configurations across All, Gumbel-$\lambda_{U}$, and strong baselines such as ReliefF and L1EN are very close, and the 95% ROC-AUC bootstrap confidence intervals overlap substantially, consistent with the expectation that differences largely reflect sampling variability (and modest model-fitting/calibration effects) rather than screening gains. Nevertheless, Gumbel-$\lambda_{U}$ paired with Random Forest achieves the numerically highest ROC-AUC (0.867; 95% CI [0.804, 0.920]), indicating that the tail-aware ranking remains competitive and clinically coherent even when dimensionality reduction is not possible.
The wider ROC-AUC confidence intervals on the Pima dataset are expected due to its smaller sample size, which increases sampling variability; conversely, the larger CDC cohort yields tighter intervals.
@@
On CDC (21 features), upper-tail filtering reduces the feature set by $\sim$52%. This introduces a minor but statistically significant performance trade-off against using the full feature set, but our method’s key advantage is highlighted by its statistically higher performance compared to other filter baselines, demonstrating its efficiency and ranking power. On PIMA (8 features), where selection is not applicable, our criterion yields clinically coherent rankings and maintains discrimination (no significant AUC differences), i.e., it does no harm in low-dimensional clinical settings. Together, these results support the method’s robustness, interpretability, and practicality across dataset types.
@@
Clarification (clinical practice vs study contribution).
The points below (for both CDC and PIMA) are not presented as new clinical guidelines. We emphasize that our study does not propose new interventions; rather, it provides a dependence-based prioritization lens from the upper-tail ranking that complements, and does not replace, established clinical practice and clinician judgment. Instead, we (i) reference established screening/management practices commonly used in diabetes and cardiometabolic care to aid interpretation, and (ii) highlight study-informed implications that arise specifically from our upper-tail dependence ranking, which is intended to prioritize attention toward the highest-risk strata rather than to propose new interventions.
@@
In this study, we developed a new supervised filter method for feature selection using the Gumbel copula’s upper-tail dependence coefficient, $\lambda_{U}$ interpreted as a rank-based tail concordance score. By focusing on how extreme values in predictors and the target label occur together, our method provides a fresh way to find features that matter most in high-risk situations. Our thorough evaluation, supports that this approach is statistically well-motivated, effective in our experiments, and computationally efficient.
We tested our method on two different diabetes datasets and found strong support for its usefulness. On the large CDC public health dataset, our Gumbel-$\lambda_{U}$ selector was competitive and was the fastest method, running much quicker than other approaches. It also reduced the number of features by approximately 52% (from 21 to 10), and on this parsimonious set it achieved statistically higher performance than standard filters such as Mutual Information and mRMR, while being statistically indistinguishable from the strong ReliefF baseline. On the PIMA dataset (8 predictors), our ranking achieved the numerically highest ROC-AUC while remaining statistically indistinguishable from strong baselines, functioning as a ranking-only clinical sanity check. Although it was not the fastest on this smaller dataset, its speed was still close to other standard filters, showing that its strong performance did not come at a high computational cost. In both cases, our method consistently highlighted predictors that are clinically plausible and consistent with prior knowledge, showing it can identify meaningful and understandable signals.
@@
Limitation.
The proposed screening procedure scores each predictor individually (purely marginally) via its upper-tail dependence with the outcome and therefore does not explicitly capture interaction effects among predictors. As a result, a feature that is weak marginally (or marginally uncorrelated) but informative through joint effects may be under-ranked or missed. This is a known limitation of marginal screening and is explicitly discussed in the SIS literature [46]. Group-based screening extensions have been proposed to better capture joint effects; for example, DC-SIS can be directly employed to screen grouped predictors (and multivariate responses) [47]. In future work, we plan to extend our framework to interaction-aware dependence screens (e.g., groupwise tail-dependence screening for small predictor subsets and conditional tail-dependence criteria) while retaining the interpretability and computational efficiency of the current filter.
@@
The idea of finding features that are extreme alongside high-risk outcomes can be used in many areas of biomedical research, especially where there is a lot of data. For example, in genomics and transcriptomics, this approach can help identify genes that are highly active in aggressive cancers. In proteomics and metabolomics, it can find molecules that are most abundant in severe disease or in response to treatment. This method can also improve clinical trial analysis by spotting unusual lab values that predict which patients might benefit most or be at higher risk, making it easier to focus on the right groups. In neuroimaging, it provides a new way to look at fMRI or PET scans by highlighting brain regions with unusually high activity linked to certain neurological disorders, adding to what we learn from average-based analyses. All of these uses aim to build more accurate or refined personalized risk models, helping to identify patients who are most at risk so they can get early and targeted care.
@@
In summary, using copula-based upper-tail dependence offers an efficient and interpretable way to select features. By prioritizing predictors associated with outcomes in the upper tail and performing as well as or better than strong baselines on both large and small datasets in our experiments, this method can be a useful tool for machine learning applications in health data. The suggested extensions above outline ways to strengthen the theory, handle real-world data challenges, and evaluate the approach on additional biomedical settings.
@@
The CDC Diabetes Health Indicators dataset has been accessed from the UCI Machine Learning Repository using the link https://doi.org/10.24432/C53919 [2]. The PIMA Indians Diabetes dataset originates from NIDDK and Smith et al. [41]; a curated copy used for this work was accessed via Kaggle at https://www.kaggle.com/datasets/uciml/pima-indians-diabetes-database.
@@
All code to reproduce the analyses is available at https://github.com/agnivibes/copula-feature-selection.
Author contributions
A.A.: Conceived and designed the study, developed the methodology, ran the experiments, analyzed the results, and wrote the manuscript. M.M.M. and S.H.: Supported the literature review and assisted in manuscript preparation. A.M.: Interpreted the results in the public health context.